OpenStudy (anonymous):

(sqrt2+5sqrt10)(sqrt2-5sqrt10) multiply

OpenStudy (karatechopper):

put in calculator?

OpenStudy (anonymous):

\[(a+b)(a-b)=a^2-b^2\] \[(\sqrt{2}+5\sqrt{10})(\sqrt{2}-5\sqrt{10})=(\sqrt{2})^2-(5\sqrt{10})^2\] \[=2-250=248\]

OpenStudy (anonymous):

Thanks I appreciate the clarification. Sometimes a fresh pair of eyes :)